# ARTICLE TYPE PAGURI: a user experience study of creative interaction with text-to-music models

### v1

@@ 4.4 Consistency, Expectations, and Quality
Table 2 reports the results related to the Model evaluation survey. Each participant was asked to take the survey after every interaction with the model. The survey aimed to gain insight into participants’ perceptions regarding three aspects: the consistency between the generated audio files and input prompts, the consistency between the generated audio samples and user expectations, and the overall quality of the audio(s) output concerning the user expectations. Only a limited part of the participants considered the overall generation to be entirely aligned with their expectations, both in terms of content and quality. The same holds for the consistency between audio and input prompts. The majority of participants rated all three aspects between and (on a Likert scale from 1 to 5), suggesting that they perceived the generated audio to align reasonably well with their expectations and the input prompts provided. Several participants were positively surprised by the audio generated, as it exceeded their expectations. However, it is important to acknowledge that the subjective nature of the expectation scale presents a limitation in this study. We did not collect pre-generated user expectation data, but we intend to address this in future experimental designs.
@@ 4.5 Integration of TTM models in the creative process
At the end of the interaction process, we asked participants to complete a final questionnaire, containing more general questions regarding the experience. Between them, there were some open and multiple-choice questions focusing on the integration of TTM models in the creative process. In Table 3 we report answers to questions evaluated on a Likert scale, while in the remainder of the section, we discuss the open questions contained in the final questionnaire.
@@
In the same questionnaire, we asked in what context the users would use the audio generated by the model both in the case of original and personalized models. The majority of the participants provided expected answers related to the use of samples/loops/beats for music production, for sound design, or as an inspiration tool. Others provided original suggestions, probably related to their personal work and creative context, such as the use for foley sounds generation, or P13*: ”during dance lectures”, and P14*: ”I would use it to build musical tracks for dance choreography”. The latter two are particularly interesting, suggesting that when considering the possible applications of TTM models, it is necessary to also consider the background of the specific user, which might desire a final applications not imagined beforehand by the designers of the models. Some participants, referring to the personalized model, suggested that ”The personalized model must adhere more closely to the specifications to be used” (P17*), or ”I would not use it because it did not meet my expectations” (P10*). This suggests that the model’s adherence to the desired objectives is important to users.
@@ 5. Discussion
High-quality generated samples are not all that matters. Sometimes they are expected by users, especially in music production, but there are contexts where they are not essential. In music creation, for instance, decent audio samples from TTM models can often suffice to stimulate creativity and be directly integrated into musicians’ workflows. Even if most of the participants did not rate the audio quality as the best (most of them rated out of the audio quality of the generated sample - Table 2), they stated that they would use the generated sounds directly, as they are, for inspiration, beat-making, music production, and even for crafting tracks for dance practice. This suggests that users which use TTM models to generate sounds for inspirational purposes may not need high-quality sounding generated audio. Instead, they might prioritize the creativity and novelty of the generated sounds over quality.
@@
Controllability and editability are at the forefront of requested features. Some limitation of the model considered in this study concerns the lack of editability, allowing the users to actively shape the direction of the audio generation process based on their choices. Participants expressed the desire to have more control on the generation process, for example, to modify only a specific section of the generated music track. It is important then, that the next generation of TTM models will also include proper user interfaces, which allow the users to set and modify specific parameters to guide the generation of the desired audio. Moreover, the graphical interfaces cannot be the same for every user but need to be contextualized according to the final users and their final tasks. In fact, some of these features have been integrated in the just released models (at the time of writing), such as Stable Audio 2.0 (from Stability AI) and Project Music GenAI Control (from Adobe).
@@ 6. Conclusions and future work
7. Ethics Statement
We acknowledge that diversity of nationality and education level of the pool of participants is limited. We however believe that their interest and knowledge and AI makes the results and their comments and suggestion interesting to the wider research community. Participants were able to use samples of their own to personalize the model. These samples might also contain music under copyright, we would like to stress that no music generated through the model was made available publicly.

### v2

@@ 4.4 Consistency, Expectations, and Quality
Fig. 3 reports the results related to the Model evaluation survey. Each participant was asked to take the survey after every interaction with the model. Each audio(s) generation corresponds to one interaction with the model. Fig. 3 reports the answers for all the different interactions of all participants. The survey aimed to gain insight into participants’ perceptions regarding three aspects: the consistency between the generated audio files and input prompts, the consistency between the generated audio samples and user expectations, and the overall quality of the audio(s) output concerning the user expectations. Only a limited part of the participants considered the overall generation to be entirely aligned with their expectations, both in terms of content and quality. The same holds for the consistency between audio and input prompts. The majority of participants rated all three aspects between and (on a Likert scale from 1 to 5), suggesting that they perceived the generated audio to align reasonably well with their expectations and the input prompts provided. Several participants were positively surprised by the audio generated, as it exceeded their expectations. However, it is important to acknowledge that the subjective nature of the expectation scale presents a limitation in this study. We did not collect pre-generated user expectation data, but we intend to address this in future experimental designs.
@@
4.4.1 Example of interaction with the model
As pointed out by the discussed results, it is hard to draw general conclusions and it is important to consider the subjectivity of the interaction with TTMs. For these reasons, we report two examples of interactions of users with PAGURI, by selecting them depending on their answer to question Q18 of the survey, namely What is your relationship with music? We select P5, who answered DJ or/and music producer, and P16, who instead answered I simply listen to music. The rationale behind this choice is to loosely analyze the impact of musical knowledge on the subjective evaluation of the interaction with PAGURI. While a plethora of in-depth granular analyses could be performed, we believe that they should be planned in the questionnaire by design, which would be out-of-scope for the presented work and thus best left for future works.
Fig. 4 reports the users’ interactions with PAGURI by showing how the answers to the questions considered in Fig. 3 evolve during the experiment. Table 1 shows the corresponding prompts used at each iteration by both considered participants. The blue bars indicate that the music generative model was not personalized (at the corresponding iteration), while the red color indicates that the model has undergone the personalization procedure. As it is possible to observe, both users start by experimenting with the model and then proceed to personalize it with samples of choice before the third iteration. It is interesting how the two users interacted with the model. While P5, supposedly more music savvy, experiments by applying effects to the personalized instrument, P16 seems to take the procedure less ”seriously”, by also experimenting with more uncoherent prompts.
@@ 4.5 Integration of TTM models in the creative process
At the end of the interaction process, we asked participants to complete a final questionnaire, containing more general questions regarding the experience. Between them, there were some open and multiple-choice questions focusing on the integration of TTM models in the creative process. In Fig. 5 we report answers to questions evaluated on a Likert scale, while in the remainder of the section, we discuss the open questions contained in the final questionnaire.
@@
In the same questionnaire, we asked in what context the users would use the audio generated by the model both in the case of original and personalized models. The majority of the participants provided expected answers related to the use of samples/loops/beats for music production, for sound design, or as an inspiration tool. Others provided original suggestions, probably related to their personal work and creative context, such as the use for foley sounds generation, or P13*: ”during dance lectures”, and P14*: ”I would use it to build musical tracks for dance choreography”. The latter two are particularly interesting, suggesting that when considering the possible applications of TTM models, it is necessary to also consider the background of the specific user, who might desire a final application not imagined beforehand by the designers of the models. Some participants, referring to the personalized model, suggested that ”The personalized model must adhere more closely to the specifications to be used” (P17*), or ”I would not use it because it did not meet my expectations” (P10*). This suggests that the model’s adherence to the desired objectives is important to users.
@@ 5. Discussion
High-quality generated samples are not all that matters Sometimes they are expected by users, especially in music production, but there are contexts where they are not essential. In music creation, for instance, decent audio samples from TTM models can often suffice to stimulate creativity and be directly integrated into musicians’ workflows. Even if most of the participants did not rate the audio quality as the best (most of them rated out of the audio quality of the generated sample - Fig. 3), they stated that they would use the generated sounds directly, as they are, for inspiration, beat-making, music production, and even for crafting tracks for dance practice. This suggests that users who use TTM models to generate sounds for inspirational purposes may not need high-quality sound-generated audio. Instead, they might prioritize the creativity and novelty of the generated sounds over quality.
@@
Controllability and editability are at the forefront of requested features. Some limitation of the model considered in this study concerns the lack of editability, allowing the users to actively shape the direction of the audio generation process based on their choices. Participants expressed the desire to have more control over the generation process, for example, to modify only a specific section of the generated music track. It is important then, that the next generation of TTM models will also include proper user interfaces, which allow the users to set and modify specific parameters to guide the generation of the desired audio. Moreover, the graphical interfaces cannot be the same for every user but need to be contextualized according to the final users and their final tasks. In fact, some of these features have been integrated into the just-released models (at the time of writing), such as Stable Audio 2.0 (from Stability AI) and Project Music GenAI Control (from Adobe).
5.1 User study limitations
As with any user study, also PAGURI suffers from limitations, some due to circumstances some due to inherent limitations of text-to-music and generative models in general.
We acknowledge that the diversity of nationality and education level of the pool of participants is limited. However, we believe that their interest and knowledge of music and AI-related tools make the results and their comments and suggestions interesting to the wider research community. Even though we tried to enlarge the number of participants by publishing the advert on several relevant mailing lists and by providing the opportunity to participate online, the time-consuming nature of the experiment made it hard to reach a wider public.
The limitations in terms of diversity can be translated also to the TTM models considered. Indeed, the music generated by the AudioLDM2 model, as well as similar models (Copet et al.,, 2024; Kreuk et al.,, 2022), frequently reflects mainstream Western music, which aligns with the cultural context in which these models were trained, and the diversity of the data used during training. These datasets, even if mostly undisclosed, are often based on Western music genres and, consequently, they do not generate music from underrepresented world cultures.